# Instructive Dialogue Summarization with Query Aggregations

### v2

@@ 3.3 Model Training
Models Params ROUGE-1 ROUGE-2 ROUGE-L BS $F_{1}$ $Pre$ $Rec$ $F_{1}$ $Pre$ $Rec$ $F_{1}$ $Pre$ $Rec$ Pointer-Generator - 40.1 - - 15.3 - - 36.6 - - - BART 400M 53.0 59.0 52.8 28.4 32.1 28.2 44.2 49.3 44.0 53.3 MV-BART 400M 53.9 55.7 57.4 28.4 29.3 30.6 44.4 45.7 47.5 53.6 Coref-BART 400M 53.7 56.9 56.4 28.5 30.5 29.7 44.3 46.9 46.5 53.5 ConDigSum 400M 54.3 56.0 57.6 29.3 30.4 31.2 45.2 46.6 48.0 54.0 GPT-3-finetune $*{}^{*}$ 53.4 - - 29.8 - - 45.9 - - - Alpaca 7B 28.2 26.0 39.8 5.7 5.1 8.3 20.5 19.2 29.0 19.4 Flan-T5-XXL 11B 52.6 62.6 50.0 28.5 34.1 27.1 44.1 52.5 41.9 53.2 Flan-UL2 20B 53.3 60.3 52.5 28.0 32.0 27.7 44.1 50.0 43.3 53.5 ChatGPT 175B 32.7 22.4 70.2 12.3 8.4 27.1 24.7 16.9 53.6 32.5 InstructDS $*{}^{*}$ 55.3 58.8 57.5 31.3 33.5 32.6 46.7 49.7 48.6 55.5 w/ reference summary length ChatGPT 175B 40.8 39.3 43.4 13.7 13.2 14.6 31.5 30.5 33.4 40.0 InstructDS $*{}^{*}$ 58.4 58.5 58.8 32.8 32.9 33.0 48.9 49.0 49.2 58.5
Table 3: ROUGE scores on SAMSum test set. The results are divided into two blocks: dedicated dialogue summarization models and general-purpose LLMs. “w/ reference summary length” indicates reference summary lengths are provided in instructions. $*{}^{*}$ indicates 37.7M trainable parameters.
@@ Acknowledgement
[…]
@@ Appendix B Templates and Examples
Case #1 from SAMSum Dialogue: Marsha: Guys, we’ve planned the trip with John last night as we promised Cynthia: great, thank you for that Marsha: but of course you have to agree on that Mohammad: sure, but I really trust you Gavin: me too Marsha: so as we decided last time, we will spend a week just on the beach Marsha: we all wanted some calm, nice place, right? Gavin: yes!!! Marsha: John found this little hotel in Nosy Be Marsha: $<$file_normal-_\__photo$>$ Marsha: and their website: $<$file_normal-_\__other$>$ Marsha: it’s quite basic the website, but it may actually be a good sign Gavin: I love it! Cynthia: it seems just perfect Marsha: I know Marsha: it’s more expensive than hotels on the mainland Marsha: but it seems to be the most beautiful place in Madagascar Mohammad: so let’s go there, it’s still cheaper than Italy for example Marsha: exactly! Summary - Human Reference Marsha and John planned the trip. They will spend a week on the beach with Cynthia, Mohammad and Gavin. They will all stay in a hotel in Nosy Be. Summary - BART Marsha, Cynthia, Mohammad and Gavin are going to Madagascar. They will spend a week on the beach. John found a hotel in Nosy Be. Summary - Alpaca The group agreed to spend their vacation on a beach in the small hotel John had found inNosiv Be, Madgascar. It was more expensivethan hoteles on main land,but it was the best place according to the group. They all agreed that it would be worth the money. Summary - Flan-UL2 Marsha, Cynthia, Mohammad and Gavin are going to Madagascar. They will spend a week on the beach in Nosy Be. Summary - ChatGPT Marsha informs the group that John had planned the trip, and they all agreed to spend a week on the beach. She shows them a picture of a hotel in Nosy Be and its website, which they all like despite it being more expensive than hotels on the mainland. Mohammad mentions that it is still cheaper than Italy and everyone agrees to go there. Summary - InstructDS Marsha, John, Cynthia, Mohammad and Gavin will spend a week on the beach in Nosy Be. It’s more expensive than hotels on the mainland, but it’s still cheaper than Italy.

### v3

@@ 3.3 Model Training
Models Params ROUGE-1 ROUGE-2 ROUGE-L BS $F_{1}$ $Pre$ $Rec$ $F_{1}$ $Pre$ $Rec$ $F_{1}$ $Pre$ $Rec$ Pointer-Generator - 40.1 - - 15.3 - - 36.6 - - - BART 400M 53.0 59.0 52.8 28.4 32.1 28.2 44.2 49.3 44.0 53.3 MV-BART 400M 53.9 55.7 57.4 28.4 29.3 30.6 44.4 45.7 47.5 53.6 Coref-BART 400M 53.7 56.9 56.4 28.5 30.5 29.7 44.3 46.9 46.5 53.5 ConDigSum 400M 54.3 56.0 57.6 29.3 30.4 31.2 45.2 46.6 48.0 54.0 GPT-3-finetune 175B 53.4 - - 29.8 - - 45.9 - - - Alpaca 7B 28.2 26.0 39.8 5.7 5.1 8.3 20.5 19.2 29.0 19.4 Flan-T5-XXL 11B 52.6 62.6 50.0 28.5 34.1 27.1 44.1 52.5 41.9 53.2 Flan-UL2 20B 53.3 60.3 52.5 28.0 32.0 27.7 44.1 50.0 43.3 53.5 ChatGPT 175B 32.7 22.4 70.2 12.3 8.4 27.1 24.7 16.9 53.6 32.5 InstructDS 3B 55.3 58.8 57.5 31.3 33.5 32.6 46.7 49.7 48.6 55.5 w/ reference summary length ChatGPT 175B 40.8 39.3 43.4 13.7 13.2 14.6 31.5 30.5 33.4 40.0 InstructDS 3B 58.4 58.5 58.8 32.8 32.9 33.0 48.9 49.0 49.2 58.5
Table 3: ROUGE scores on SAMSum test set. The results are divided into two blocks: dedicated dialogue summarization models and general-purpose LLMs. “w/ reference summary length” indicates reference summary lengths are provided in instructions. indicates 37.7M trainable parameters.
@@ Acknowledgement
This research was supported by funding from the Institute for Infocomm Research (I2R), A*STAR, Singapore, and DesCartes: The National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) program. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore.
@@ Appendix B Templates and Examples
Case #1 from SAMSum Dialogue: Marsha: Guys, we’ve planned the trip with John last night as we promised Cynthia: great, thank you for that Marsha: but of course you have to agree on that Mohammad: sure, but I really trust you Gavin: me too Marsha: so as we decided last time, we will spend a week just on the beach Marsha: we all wanted some calm, nice place, right? Gavin: yes!!! Marsha: John found this little hotel in Nosy Be Marsha: $<$file__\__photo$>$ Marsha: and their website: $<$file__\__other$>$ Marsha: it’s quite basic the website, but it may actually be a good sign Gavin: I love it! Cynthia: it seems just perfect Marsha: I know Marsha: it’s more expensive than hotels on the mainland Marsha: but it seems to be the most beautiful place in Madagascar Mohammad: so let’s go there, it’s still cheaper than Italy for example Marsha: exactly! Summary - Human Reference Marsha and John planned the trip. They will spend a week on the beach with Cynthia, Mohammad and Gavin. They will all stay in a hotel in Nosy Be. Summary - BART Marsha, Cynthia, Mohammad and Gavin are going to Madagascar. They will spend a week on the beach. John found a hotel in Nosy Be. Summary - Alpaca The group agreed to spend their vacation on a beach in the small hotel John had found inNosiv Be, Madgascar. It was more expensivethan hoteles on main land,but it was the best place according to the group. They all agreed that it would be worth the money. Summary - Flan-UL2 Marsha, Cynthia, Mohammad and Gavin are going to Madagascar. They will spend a week on the beach in Nosy Be. Summary - ChatGPT Marsha informs the group that John had planned the trip, and they all agreed to spend a week on the beach. She shows them a picture of a hotel in Nosy Be and its website, which they all like despite it being more expensive than hotels on the mainland. Mohammad mentions that it is still cheaper than Italy and everyone agrees to go there. Summary - InstructDS Marsha, John, Cynthia, Mohammad and Gavin will spend a week on the beach in Nosy Be. It’s more expensive than hotels on the mainland, but it’s still cheaper than Italy.